# LocationReasoner: Evaluating LLMs on Real-World Site Selection Reasoning

### v1

@@ 3.1 Main Results
simple 89.81% 71.30% 0.66 0.61 0.60
medium 85.42% 55.21% 0.57 0.46 0.45
hard 88.24% 37.25% 0.26 0.17 0.15
OpenAI 4o overall 87.91% 55.00% 0.50 0.42 0.40
simple 91.82% 81.82% 0.74 0.74 0.72
medium 89.58% 73.96% 0.67 0.64 0.61
hard 88.24% 53.92% 0.42 0.40 0.38
OpenAI o4-mini overall 89.94% 69.99% 0.61 0.59 0.57
@@
simple 66.67% 35.19% 0.39 0.34 0.31
medium 47.92% 21.88% 0.14 0.08 0.07
hard 30.39% 9.80% 0.03 0.01 0.01
Claude 3 Haiku overall 48.69% 23.00% 0.19 0.15 0.13
simple 67.59% 52.78% 0.51 0.50 0.48
medium 65.62% 54.17% 0.43 0.40 0.39
hard 46.08% 26.47% 0.22 0.21 0.18
Claude 3.5 Haiku overall 59.80% 44.00% 0.39 0.37 0.35
Agentic Workflow
@@
Limited overall performance. The overall model performance on the LocationReasoner benchmark remains limited, as most models fail to generate complete and accurate outputs across varying difficulty levels. The average perfect pass rate across 8 different LLMs is only 49.61%. Even the best-performing model, OpenAI o4-mini, achieves only a 69.99% perfect pass rate and an overall F1 score of 0.57. This reflects a significant gap between current LLM capabilities and the demands of real-world complex reasoning tasks.
@@
Difficulty sensitivity. Across all models, performance consistently degrades as task difficulty increases. All key metrics decline from simple to medium to hard queries. Specifically, the average perfect pass rate on hard queries is only 33.82%, which is much lower than 61.49% of easy queries. This trend highlights that while some LLMs may be capable of handling basic filtering or single-constraint tasks, they struggle to generalize to multi-step reasoning and more compositional decision-making challenges.
@@
Reasoning LLMs offer significant improvements. Despite the DeepSeek family, we observe significant progress in all other three families made by their reasoning models compared with their baseline counterparts. For example, Gemini 2.5, advertised as a reasoning-optimized model, achieves an overall F1 score of 0.48 and a perfect pass rate of 62%. Its predecessor, Gemini 1.5, reaches only a 27% perfect pass rate and an F1 score of 0.21. However, the gains remain far from transformative. Gemini 2.5 still fails over 35% of simple queries and more than half of medium ones. This underscores that even when reasoning-optimized models do improve over prior versions, their performance still falls short of the reliability required for real-world reasoning tasks.
@@
Agentic strategies do not guarantee better results. Agentic strategies such as ReAct and Reflexion do not outperform direct code generation. Despite being designed to simulate step-by-step human reasoning, ReAct + Reflexion achieves only a 33% perfect pass rate and an overall F1 score of 0.24, which is substantially lower than the best-performing code generation LLMs. This suggests that agentic stratety alone is insufficient without more reliable reasoning abilities of the base model. We provide detailed explanations on the inferior performance of agentic strategies in Appendix A.
@@ 4 Error Analysis
Agents consistently struggle with five primary failure types, each corresponding to the categories shown in Figure 2. These errors often overlap, but their root causes reflect distinct reasoning weaknesses that hinder constraint satisfaction and plan validity.
@@
  • Logic Error. Logic errors encompass broken reasoning chains, incorrect constraint handling, and improperly ordered filtering logic. This includes sequential filtering mistakes where each constraint is applied too early, altering the input for future conditions. In ReAct, logic errors often emerge as inconsistent multi-step chains, where prior outputs are forgotten or overwritten. In code generation, they tend to surface as incorrect condition checks or missing final evaluations.
  • Edge Case Error. LLMs frequently misinterpret constraints that require careful logic around null values or boundary conditions. A common example involves requests like “fewer than 3 competitors,” where agents exclude zones with zero competitors–despite these being valid. These failures stem from inflexible logic structures and an inability to reason inclusively.
  • Tool Error. This category captures failures where the model misuses, miscalls, or entirely ignores in-house tools. Examples include passing incorrect argument formats, selecting a tool misaligned with the constraint type, or chaining incompatible tool outputs. Code generation sometimes bypass the toolset and attempt to write their own versions of utility functions, which introduces further bugs and inconsistencies. These issues reveal brittle tool affordance understanding and poor mapping between language and tool usage.
@@
We evaluate whether reasoning models can correct the errors made by their nonreasoning counterparts. Table 3 shows the results where Correction Rate is defined as the proportion of test cases in which the reasoning model produces a correct output while the non-reasoning model fails. Our results indicate that reasoning models are particularly effective at correcting failures on medium queries. This is likely because medium queries contain a high concentration of logic-related errors, which are well-suited to correction through enhanced reasoning. In contrast, easy queries tend to involve edge cases, while hard queries are often dominated by syntax issues and execution errors, where reasoning alone is insufficient to guarantee success. For example, OpenAI’s o4-mini corrects 25.36% more errors over GPT-4o in medium queries than hard or simple ones. However, the extent of improvement varies considerably across model families, highlighting that the effectiveness of reasoning depends not only on the presence of reasoning mechanisms but also on how well they are integrated and executed. We also conducted prompt ablation experiment and the results are provided in Appendix B.
@@ 5 Case Study
Figure 3: Example test case failures: over-reasoning (left) and sequential filtering (right).
  • Over-reasoning (Code and ReAct). Agents frequently demonstrate over-reasoning by performing unnecessary operations that go beyond the requirements of the prompt. This often stems from misinterpreting implicit cues. In direct code generation, agents tend to apply extra filters that overly constrain the result set, introducing conditions not specified in the query. In the ReAct setting, the model continues to invoke tools and apply filters even after all constraints have been met, lacking a clear signal for termination. As illustrated in Figure 3(left), the ReAct agent first retrieves parking capacity correctly, then unnecessarily proceeds to filter for the presence of shopping malls and many other unrelated tools. The agent accumulates irrelevant actions until it reaches the step limit and fails to return an answer, despite having the correct information early on.
  • Sequential Filtering (Code). In complex queries involving multiple chains of constraints, the model often applies filters sequentially rather than evaluating the logic holistically. As shown in Figure 3(right), the query contains an OR condition between competitor count and population. The correct approach (left) evaluates both branches independently before merging the results. However, the LLM agent (right) applies filters sequentially, first by competitor count and then by population, prematurely discarding zones that could have satisfied the second condition. This flawed execution leads to valid zones being excluded, revealing a logical failure where sequential filtering overrides the intended logical structure.
@@ 6 Related Work
LLM Reasoning. Reasoning represents a key capability of advanced LLMs, enabling them to decompose tasks into smaller subproblems, perform structured search, and make decisions for complex problem-solving. Various techniques have been proposed to enhance the reasoning abilities of LLMs. Prompting methods wei2022chain ; yao2023tree ; besta2024graph , such as chain-of-thought (CoT) prompting wei2022chain , encourage models to generate intermediate and structured reasoning steps before producing final answers. These sequential reasoning patterns can be further extended into more expressive structures, such as trees yao2023tree , for handling more complex reasoning workflows. Beyond prompting, recent research has shown that reasoning capabilities can also be distilled into LLMs’ generation itself through post-training strategies guo2025deepseek ; snell2024scaling ; brown2024large ; zelikman2024quiet . In particular, reinforcement learning is often used to fine-tune models to produce high-quality reasoning trajectories, such as CoT-style paths guo2025deepseek ; havrilla2024teaching ; kumar2024training ; carta2023grounding . In parallel, agentic approaches offer an alternative paradigm, equipping LLMs with tools, external memory, and planning mechanisms to solve complex tasks wiesinger2024agents ; yao2023react ; shinn2023reflexion ; zhao2024expel ; shen2023hugginggpt ; schick2023toolformer . For instance, the ReAct framework yao2023react enables in-context learning by prompting models to engage in iterative thought-action-observation cycles. In this work, we systematically evaluate both LLMs and reasoning-enhanced strategies on practical real-world reasoning tasks through our proposed benchmark.
@@
Benchmarks for LLMs and Agents. Benchmarks play an essential role in the development and evaluation of LLMs. Fundamental language capabilities are typically assessed through general-purpose benchmarks targeting language understanding and factual question answering srivastava2022beyond ; wang2024mmlu ; wang2024mmlu ; rein2024gpqa . In addition, domain-specific benchmarks have been introduced to evaluate LLMs’ proficiency in specialized areas such as biology chen2023bioinfo and law fei2023lawbench . As reasoning has emerged as a central focus in recent LLM research, benchmarks tailored specifically to reasoning have become increasingly important. Given the complexity of evaluating LLM reasoning, existing benchmarks often focus on domains with well-defined verifiers, where checking the correctness of a solution is significantly easier than generating it. Two such domains–mathematics and programming–have become prominent battlegrounds for reasoning-oriented LLMs math500 ; quan2025codeelo ; aime2024 , exemplified by benchmarks such as MATH500 math500 and CodeForces quan2025codeelo . However, constructing benchmarks for practical reasoning tasks remains challenging, largely due to the difficulty of defining objective metrics and the high cost of manual annotation xie2024travelplanner . In contrast, our benchmark evaluates LLMs and agents on real-world site selection reasoning while maintaining full scalability through automatic verification, requiring no human annotation. This enables robust, reproducible, and extensible evaluation of reasoning capabilities in realistic decision-making contexts.
@@ 7 Conclusion
In this paper, we present LocationReasoner, a benchmark designed to evaluate LLMs and agents on real-world site selection reasoning tasks, supported by a sandbox environment enriched with built-in tools and curated datasets. Through systematic evaluation of both general-purpose and reasoning-oriented LLMs from multiple providers, along with various agentic strategies, we find that current models perform poorly–with even the most advanced reasoning LLM failing to solve over 30% of the queries. Our analysis identifies key bottlenecks, particularly in over-reasoning and the challenges of sequential problem-solving, which limit the effectiveness of LLMs in complex real-world reasoning scenarios. LocationReasoner serves as a scalable and extensible testbed: new constraints can be added and automatically validated without the need for human annotation. We hope LocationReasoner will drive progress in enhancing the reasoning capabilities of LLMs and contribute to the advancement of general intelligence for solving practical, high-stakes problems.
Appendix A Reasoning Architecture Comparison
@@
Appendix B Prompt ablation

### v2

@@ 3 Benchmarking Results
[…]
@@
[…]
@@
Table 3 summarize the results of 316 site selection queries (114 simple, 102 medium, 102 hard) using Boston data, from which we have the following observations.
@@
Limited overall performance. The overall model performance on the LocationReasoner benchmark remains limited, as most models fail to generate complete and accurate outputs across varying difficulty levels. The average perfect pass rate for 8 different LLMs is only 49.61%. Even the best-performing model, OpenAI o4-mini, achieves only a 69.99% perfect pass rate and an overall F1 score of 0.57. This reflects a significant gap between current LLM capabilities and the demands of complex real world reasoning tasks.
@@
Difficulty sensitivity. Across all models, performance consistently degrades as task difficulty increases. All key metrics decline from simple to medium to hard queries. Specifically, the average perfect pass rate on hard queries is only 33.82%, which is much lower than 61.49% of simple queries. This trend highlights that while some LLMs may be capable of handling basic filtering or single-constraint tasks, they struggle to generalize to multi-step reasoning and more compositional decision-making challenges.
@@
Reasoning LLMs offer significant improvements. With the exception of the DeepSeek family, the other three families show substantial improvements in their reasoning models compared to their non-reasoning baseline counterparts. For example, Gemini 2.5, advertised as a reasoning-optimized model, achieves an overall F1 score of 0.48 and a perfect pass rate of 62%. Its predecessor, Gemini 1.5, reaches only a 27% perfect pass rate and an F1 score of 0.21. However, the gains remain far from transformative. Gemini 2.5 still fails over 35% of simple queries and more than half of medium ones. This underscores that even when reasoning-optimized models do improve over their previous versions, their performance still falls short of the reliability required for real-world reasoning tasks.
@@
Agentic strategies do not guarantee better results. Agentic strategies such as ReAct and Reflexion do not outperform direct prompting. Despite being designed to simulate step-by-step human reasoning, ReAct + Reflexion achieves only a 33% perfect pass rate and an overall F1 score of 0.24, which is substantially lower than the best-performing LLMs using direct prompting. This suggests that agentic strategy alone is insufficient without more reliable reasoning abilities of the base model. We provide detailed explanations on the inferior performance of agentic strategies in Appendix A.1.
4 Scalability and Statistical Robustness
The automated query generation and verification property of LocationReasoner ensures its scalability to incorporate more queries and more datasets. In this section, we re-ran the experiments in two additional cities: New York and Tampa. These cities feature distinctive urban layouts and provide a more comprehensive testbed for spatial reasoning. From Table 4 in Appendix we observe that the results on these new datasets are consistent with our initial findings, showing that even top-performing models struggle with complex spatial reasoning, and highlighting the persistent challenges in this domain. Particularly, the best-performing model, Gemini 2.5, can only achieve a Perfect Pass Rate of 58.40% on hard queries in New York, and 61.74% in Tampa. This expansion demonstrates that our framework is robust and applicable to different site selection environments.
Figure 2: Model performance convergence with increasing query count.
To evaluate the robustness of our benchmark with respect to dataset size, we further conducted an analysis tracking the Perfect Pass Rate of models as the number of queries increases using the Boston dataset. As shown in Figure 2, the performance for all models begins to converge around 150 queries, with subsequent fluctuations remaining minimal. This indicates that our choice of $\sim$300 queries per city is more than sufficient to provide a stable and statistically significant signal of model capability. Furthermore, since our queries and the sandbox environment support fully automated evaluation, the size of the benchmark can be easily and efficiently increased in future work to test even more fine-grained aspects of LLM reasoning. This focus on automated, low-cost scalability makes LocationReasoner a uniquely practical and sustainable tool for rigorously testing complex, constraint-based spatial and logical reasoning that is complementary to the tested abilities on other benchmarks.
@@ 5 Error Analysis
LLMs consistently struggle with five primary failure types, and we show the breakdown of these failures in Figure 3. These errors often overlap, but their root causes reflect distinct reasoning weaknesses that hinder constraint satisfaction and plan validity.
Logic Error encompasses broken reasoning chains, incorrect constraint handling, and improperly ordered filtering logic. This includes sequential filtering mistakes where each constraint is applied too early, altering the input for future conditions. In ReAct, logic errors often emerge as inconsistent multi-step chains, where prior outputs are forgotten or overwritten. In direct prompting, they tend to surface as incorrect condition checks or missing final evaluations.
Edge Case Error indicates that LLMs frequently misinterpret constraints that require careful logic around null values or boundary conditions. A common example involves requests like “fewer than 3 competitors,” where agents exclude zones with zero competitors–despite these being valid. These failures stem from inflexible logic structures and an inability to reason inclusively.
Tool Error captures failures where the model misuses, miscalls, or entirely ignores in-house tools. Examples include passing incorrect argument formats, selecting a tool misaligned with the constraint type, or chaining incompatible tool outputs. LLMs sometimes bypass the toolset and attempt to write their own versions of utility functions, which introduce further bugs and inconsistencies. These issues reveal brittle tool affordance understanding and poor mapping between language and tool usage.
@@
We evaluate whether reasoning models can correct the errors made by their non-reasoning counterparts. Table 3 shows the results where Correction Rate is defined as the proportion of test cases in which the reasoning model produces a correct output while the non-reasoning model fails. Our results indicate that reasoning models are particularly effective in correcting failures in medium queries. This is likely because medium queries contain a high concentration of logic-related errors, which are well suited to correction through enhanced reasoning. In contrast, easy queries tend to involve edge cases, while hard queries are often dominated by syntax issues and execution errors, where reasoning alone is insufficient to guarantee success. For example, OpenAI’s o4-mini corrects 25.36% more errors over GPT-4o in medium queries than hard or simple ones. However, the extent of improvement varies considerably across model families, highlighting that the effectiveness of reasoning depends not only on the presence of reasoning mechanisms but also on how well they are integrated and executed. We provide concrete failure examples including the generated codes as well as prompt ablation experiments in Appendix A.2-A.3.
@@ 6 Case Study
Over-reasoning (Direct Prompting and ReAct). Agents frequently demonstrate over-reasoning by performing unnecessary operations that go beyond the requirements of the prompt. This often stems from misinterpreting implicit cues. In direct prompting, agents tend to apply extra filters that overly constrain the result set, introducing conditions not specified in the query. In the ReAct setting, the model continues to invoke tools and apply filters even after all constraints have been met, lacking a clear signal for termination. As illustrated in Figure 4, the ReAct agent first retrieves parking capacity correctly, then unnecessarily proceeds to filter for the presence of shopping malls and many other unrelated tools. The agent accumulates irrelevant actions until it reaches the step limit and fails to return an answer, despite having the correct information early on.
Figure 5: Sequential Filtering
Sequential Filtering (Direct Prompting). In complex queries involving multiple chains of constraints, the model often applies filters sequentially rather than evaluating the logic holistically. As shown in Figure 5, the query contains an OR condition between competitor count and population. The correct approach (right) evaluates both branches independently before merging the results. However, the LLM agent (left) applies filters sequentially, first by competitor count and then by population, prematurely discarding zones that could have satisfied the second condition. This flawed execution leads to valid zones being excluded, revealing a logical failure where sequential filtering overrides the intended logical structure.
@@ 7 Related Work
LLM Reasoning. Reasoning represents a key capability of advanced LLMs, enabling them to decompose tasks into smaller subproblems, perform structured search, and make decisions for complex problem solving. Various techniques have been proposed to enhance the reasoning abilities of LLMs. Prompting methods (Wei et al., 2022b; Yao et al., 2023a; Besta et al., 2024), such as chain-of-thought (CoT) prompting (Wei et al., 2022b), encourage models to generate intermediate and structured reasoning steps before producing final answers. These sequential reasoning patterns can be further extended into more expressive structures, such as trees (Yao et al., 2023a), for handling more complex reasoning workflows. Beyond prompting, recent research has shown that reasoning capabilities can also be distilled into LLMs’ generation itself through post-training strategies (Guo et al., 2025; Snell et al., 2024; Brown et al., 2024; Zelikman et al., 2024). In particular, reinforcement learning is often used to fine-tune models to produce high-quality reasoning trajectories, such as CoT-style paths (Guo et al., 2025; Havrilla et al., 2024; Kumar et al., 2024; Carta et al., 2023). In parallel, agentic approaches offer an alternative paradigm, equipping LLMs with tools, external memory, and planning mechanisms to solve complex tasks (Wiesinger et al., 2024; Yao et al., 2023b; Shinn et al., 2023; Zhao et al., 2024; Shen et al., 2023; Schick et al., 2023). For instance, the ReAct framework (Yao et al., 2023b) enables in-context learning by prompting models to engage in iterative thought-action-observation cycles. In this work, we systematically evaluate both LLMs and reasoning-enhanced strategies on practical real-world reasoning tasks through our proposed benchmark.
@@
Benchmarks for LLMs and Agents. Benchmarks play an essential role in the development and evaluation of LLMs. Fundamental language capabilities are typically assessed through general-purpose benchmarks targeting language understanding and factual question answering (Srivastava et al., 2022; Wang et al., 2024b; b; Rein et al., 2024). In addition, domain-specific benchmarks have been introduced to evaluate LLMs’ proficiency in specialized areas such as biology (Chen & Deng, 2023) and law (Fei et al., 2023). As reasoning has emerged as a central focus in recent LLM research, benchmarks tailored specifically to reasoning have become increasingly important. Given the complexity of evaluating LLM reasoning, existing benchmarks often focus on domains with well-defined verifiers, where checking the correctness of a solution is significantly easier than generating it. Two such domains–mathematics and programming–have become prominent battlegrounds for reasoning-oriented LLMs (Hendrycks et al., 2021b; Quan et al., 2025; MAA, 2024), exemplified by benchmarks such as MATH500 (Hendrycks et al., 2021b) and CodeForces (Quan et al., 2025). However, constructing benchmarks for practical reasoning tasks remains challenging, largely due to the difficulty of defining objective metrics and the high cost of manual annotation (Xie et al., 2024). In contrast, our benchmark evaluates LLMs and agents on real-world site selection reasoning while maintaining full scalability through automatic verification, requiring no human annotation. This enables robust, reproducible, and extensible evaluation of reasoning capabilities in realistic decision-making contexts.
@@ 8 Conclusion
In this paper, we present LocationReasoner, a benchmark designed to evaluate LLMs and agents on real-world site selection reasoning tasks, supported by a sandbox environment enriched with built-in tools and curated datasets. Through systematic evaluation of both general-purpose and reasoning-oriented LLMs from multiple providers, along with various agentic strategies, we find that current models perform poorly, with even the most advanced reasoning LLM failing to solve over 30% of the queries. Our analysis identifies key bottlenecks, particularly in over-reasoning and the challenges of sequential problem-solving, which limit the effectiveness of LLMs in complex real-world reasoning scenarios. LocationReasoner serves as a scalable and extensible testbed: new constraints can be added and automatically validated without the need for human annotation. We hope LocationReasoner will drive progress in enhancing the reasoning capabilities of LLMs and contribute to the advancement of general intelligence for solving practical, high-stakes problems.
Appendix A Appendix
A.1 Reasoning Architecture Comparison
@@
A.2 Model Errors
These examples illustrate how LLMs perform on individual questions and the types of reasoning errors that can occur, even with advanced models.
Case 1: Edge Case — Zones with 0 Competitors Are Excluded
User Query: “I want to open a clothing store in White Plains, with the top category being Other Schools and Instruction and sub-category Exam Preparation and Tutoring. Show me zones with less than 3 competitors in the same category.”
Analysis: Claude3Haiku filters only among existing competitors, excluding zones with 0 competitors. DeepseekV3 ensures all zones are evaluated by initializing the count to 0.
Incorrect – Claude3Haiku
category_filtered = filter_pois_by_sub_category(
poi_spend_df, "Other Schools and Instruction")
competitor_counts = category_filtered[’zone_id’].value_counts().reset_index()
competitor_counts.columns = [’zone_id’, ’num_competitors’]
valid_zones = competitor_counts[
competitor_counts[’num_competitors’] < 3][’zone_id’].tolist()
Correct – Deepseek V3
all_zones = poi_spend_df[’zone_id’].unique()
category_filtered = filter_pois_by_sub_category(
poi_spend_df, "Other Schools and Instruction")
competitor_counts = category_filtered[’zone_id’].value_counts().to_dict()
zone_to_count = {zone_id: competitor_counts.get(zone_id, 0)
for zone_id in all_zones}
valid_zones = [zone_id for zone_id, count in zone_to_count.items() if count < 3]
Case 2: Prompt Misinterpretation
User Query: “I want to open a new restaurant, but I need a location with at least 50 parking spots nearby.”
Analysis: Gemini 1.5 misinterprets “50 parking spots” as requiring 50 different lots, while OpenAI 4o correctly interprets the query as needing one lot with sufficient capacity.
Incorrect – Gemini 1.5
if get_num_parking(parking_df) >= 50:
valid_zones.append(zone)
Correct – OpenAI 4o
if get_largest_parking_capacity(parking_df) >= 50:
valid_zones.append(zone)
Case 3: Code Error
User Query: “Looking to build a spa — find me areas where sub-category Advertising Agencies dominates at least 40% of 2024 spend or has 2+ parking spots.”
Analysis: Deepseek R1 crashes due to lack of error handling in low-data zones, causing a divide-by-zero error. GPT-4o correctly calls the tool and includes a guard clause to ensure stability.
Incorrect – Deepseek R1
ad_pois = filter_pois_by_sub_category(zone_pois, "Advertising Agencies")
ad_spend = get_spendparam_years(ad_pois, "RAW_TOTAL_SPEND", 2024)
total_spend = 0
if ad_spend / total_spend >= 0.4 or parking_count >= 2:
valid_zones.add(zone_id)
Correct – GPT-4o
ad_pois = filter_pois_by_sub_category(zone_pois, "Advertising Agencies")
ad_spend = get_spendparam_years(ad_pois, "RAW_TOTAL_SPEND", 2024)
total_spend = get_total_spend(zone_id, 2024)
if (total_spend > 0 and ad_spend / total_spend >= 0.4) or parking_count >= 2:
valid_zones.add(zone_id)
These examples illustrate how LLMs perform on individual questions and the types of reasoning errors that can occur, even with advanced LLMs. We have added these case studies to the revised paper to provide more concrete examples of model performance and the challenges faced in this domain.
Table 4: Performance metrics on LocationReasoner on New York and Tampa
Location Model Difficulty Delivery Rate Perfect Pass Precision Recall F1 Score
New York Gemini 1.5 Simple 55.40% 46.87% 0.35 0.28 0.31
Medium 49.24% 34.49% 0.22 0.20 0.21
Hard 38.97% 20.13% 0.10 0.07 0.08
Gemini 2.5 Simple 96.73% 78.88% 0.67 0.63 0.65
Medium 89.02% 73.61% 0.53 0.47 0.50
Hard 73.62% 58.40% 0.36 0.32 0.34
ReAct Simple 87.16% 45.54% 0.48 0.38 0.42
Medium 81.05% 33.72% 0.41 0.19 0.26
Hard 78.24% 15.15% 0.28 0.08 0.12
ReAct + Reflexion Simple 98.97% 46.28% 0.58 0.32 0.41
Medium 93.41% 29.03% 0.55 0.27 0.36
Hard 92.37% 23.95% 0.35 0.14 0.20
Tampa Gemini 1.5 Simple 50.48% 39.12% 0.29 0.27 0.20
Medium 53.76% 44.07% 0.21 0.18 0.19
Hard 34.66% 23.82% 0.12 0.06 0.08
Gemini 2.5 Simple 90.21% 71.26% 0.62 0.66 0.64
Medium 84.85% 67.89% 0.50 0.48 0.49
Hard 77.03% 61.74% 0.31 0.37 0.34
ReAct Simple 68.45% 46.78% 0.42 0.39 0.40
Medium 76.12% 33.25% 0.33 0.21 0.26
Hard 59.88% 18.92% 0.20 0.09 0.12
ReAct + Reflexion Simple 96.88% 35.62% 0.50 0.30 0.37
Medium 91.23% 39.05% 0.46 0.26 0.33
Hard 89.47% 26.73% 0.28 0.12 0.16
A.3 Prompt ablation